OpenStudy (anonymous):

$4000 for 6 years at 7% compounded monthly i=prt

OpenStudy (anonymous):

compounded intrst

OpenStudy (anonymous):

6 years = how many months

OpenStudy (anonymous):

Jordan u there

OpenStudy (anonymous):

72

OpenStudy (anonymous):

great now the formula for compounding is ???

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Amount = Principal (1 + r)^n

OpenStudy (anonymous):

\[4000\times (1+\frac{.07}{12})^{12\times 6}\]

OpenStudy (anonymous):

Amount = $ 4000 * (1+0.07)^72

OpenStudy (anonymous):

the 12 in the denominator because you get 1/12 th of your interest per month, the \(12\times 6\) in the exponent because there are 12 months in a year, times 6 years gives you the number of compounding periods

OpenStudy (anonymous):

@sheg you just gave yourself 7% per month, not per year!

OpenStudy (anonymous):

@satellite73 where it is mentioned in terms of 7% per year

OpenStudy (anonymous):

a=4142.06 b=6080.42 c=63022.28 d=4864.34

OpenStudy (anonymous):

jordan in this case 7 % PA would be their and according to it answer would be b